OpenStudy (elonasushchik):

At its closest point, Mercury is approximately 46 million kilometers from the sun. What is this distance in AU? 0.3 AU<-------- 2.4 AU 30 AU 150 million AU

OpenStudy (elonasushchik):

@Phebe

OpenStudy (elonasushchik):

@welshfella @nincompoop @acxbox22 can u guys verify my answer??

OpenStudy (acxbox22):

how many kilometers are there in 1 amu?

OpenStudy (phebe):

@dan815

OpenStudy (phebe):

@Nnesha

OpenStudy (welshfella):

is 1 astronomical unit = the distance of the earth form the sun? I seem to remember this from way back

OpenStudy (welshfella):

but i might be completely wrong ..

OpenStudy (elonasushchik):

so im understanding that my answer was wrong

OpenStudy (elonasushchik):

1 astronomical unit 1 AU = 1.4960 * 10^11 meters it is the average distance between earth and sun mercury to sun distance is = 46,000,000 * 1000 meters = 4.6 * 10^9 meters = 4.6 * 10^9 / 1.4960 * 10^11 AU = 3.0.74 / 100 = 0.0374 AU

OpenStudy (welshfella):

if thats the case the its 46 / 149.8 am

OpenStudy (elonasushchik):

the answer is C then

OpenStudy (welshfella):

distance of sun from the earth = 92.9 million miles = 149.8 million kilometres

OpenStudy (elonasushchik):

am i correct?

OpenStudy (welshfella):

no 46/149.8 = 0.30

OpenStudy (welshfella):

yes you are correct

OpenStudy (elonasushchik):

thank you @welshfella so much!<3